OpenStudy (anonymous):

olve 4x2 = 12x + 40.

OpenStudy (bahrom7893):

Try this one yourself, come on!

OpenStudy (anonymous):

solve 4^2

OpenStudy (bahrom7893):

first step should be moving everything over to one side and dividing everything by 4

OpenStudy (anonymous):

Thats the Point, I dont know how.

OpenStudy (bahrom7893):

Ok, i'll show u this ex. Try the next one: 4x^2 = 12x + 40 4x^2 - 12x - 40 = 0

OpenStudy (bahrom7893):

now divide everything by 4: x^2 - 3x - 10 = 0

OpenStudy (bahrom7893):

Now use the discriminant formula: D = (-3)^2 - 4*(1)*(-10) = 9 + 40 = 49

OpenStudy (bahrom7893):

x1 = (-b - sqrt(D))/(2a) x2 = (-b +sqrt(D))/(2a)

OpenStudy (bahrom7893):

x1 = (3 - 7)/2 = -2 x2 = (3+7)/2 = 5